OpenStudy (anonymous):

determine the radius of the circle with the given area - A=25pi

OpenStudy (anonymous):

A is pi times r squared. sooo square root of 25 is 5

OpenStudy (rulnick):

A = pi r^2 = pi * 25 so r^2 = 25 r = 5 or -5 radius cannot be negative so r = 5 done.